john is presently 12 years older than fred. four years ago john was twice as old as fred. how old is each now?
Answer by london maths tutor(243)

Let's john age be "J" and Fred age be "F".
john is presently 12 years older than fred : J = F + 12 .. equation 1
four years ago john was twice as old as fred: (J-4) = 2(F-4).. equation 2
Put equation 1 into equation 2:
(F+12)-4 = 2(F-4)
F + 12 - 4 = 2F - 8
F + 8 = 2F - 8
8 + 8 = 2F - F
16 = F
J = 16 + 12 = 28
John is 28 and Fred is 16 now.